Three carts of masses m1 = 3.50 kg, m2 = 8.00 kg, and m3 = 3.00 kg move on a frictionless, horizontal track with speeds of v1 = 7.00 m/s to the right, v2 = 3.00 m/s to the right, and v3 = 5.00 m/s to the left, as shown below. Velcro couplers make the carts stick together after colliding. (a) Find the final velocity of the train of three carts.
Give me the magnitude in m/s.
Answer #1
conservation of momentum
right direction being positive
m1v1+m2v2+m3v3=(m1+m2+m3)Vf
(3.5*7)+(8*3)+(3*-5)=14.5vf
vf=33.5/14.5=2.31m/s to the right